If k is a positive even integer , then (k+1)(k+2) could equal which of the following ?
(A) 10 (B) 20 (C) 30 (D) 40 (E) 50

Answer by richwmiller(17219)

4 * 5 = 20 but 3 is not even
What two consecutive numbers equals any of the options? 5 and 6
If k were 4 then 5*6=30
(c) 30